# SR-CACO-2: A Dataset for Confocal Fluorescence Microscopy Image Super-Resolution

### v1

@@ 8 Conclusion
The limited existing datasets for SISR microscopy are mostly private which hinders research progress. Our work comes to fill in this gap. In particular, we propose a new dataset, SR-CACO-2, for Confocal Fluorescence Microscopy Image Super-Resolution. It contains 3 scales (X2, X4, and X8) with three protein markers in different lighting conditions and about 9k real pairs of LR/HR patches. The capturing procedure of the data and the experimental protocol for SISR are straightforward and reproducible. Our benchmarking of state-of-the-art SISR methods showed that they cannot accurately produce the HR images, making this new dataset extremely challenging. Aside from the evaluation of the quality of the produced images, we conducted further analysis to assess their efficiency in downstream biology task such as cell segmentation. Several methods achieved promising results compared to LR and HR images. This dataset can contribute in driving progress in designing SISR models for microscopy. This is not limited to fixed-imaging but it extends to live-imaging. Trained models with our data can be used for live-imaging videos captured with low quality under reduced light exposure to preserve cells. This leads to fast imaging which allows the observation of instantaneous inter-cellular events with less damage to the cells. Accurately upscaling these videos with these SISR methods opens the doors for the ultimate goal which is to perform biological tasks such as cell tracking, counting, and segmentation.
@@ Acknowledgments
Appendix A Overview of Dataset
@@ A.1 Dataset statistics
We provide a nutrition label for our SR-CACO-2 dataset (Fig.12). Furthermore, Tab.6 presents more statistics about ROI distribution per HR tile for SR-CACO-2 dataset. Note that cells cover only small part of the tile in general.
@@ A.2 Data Format and Organization
where ID $\in$ [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, ’7’, ’8’, ’9’, ’10’, ’11’, ’12’, ’13’, ’14’, ’15’, ’16’, ’17’, ’18’, ’19’, ’20’, ’21’, ’22’]. Note that a tile contains all the three markers, one in each plane: plane 0 for CELL0, plane 1 for CELL1, plane 2 for CELL2. The names of the patches follows this format:

### v2

@@ 5 Conclusion
The few existing datasets for SISR microscopy are mostly private which limits research advancements. Our work seeks to address this gap. In particular, we propose a new dataset, SR-CACO-2, for Confocal Fluorescence Microscopy Image Super-Resolution. It contains 3 scales (X2, X4, and X8, in addition to HR) with three protein markers in different lighting conditions and about 9k real pairs of LR/HR patches. The procedure for data capture and the experimental protocol for the evaluation of SISR methods follow standard practices, in addition to being reproducible. Our benchmarking of state-of-the-art SISR methods indicates that they cannot accurately produce the HR images, making this new dataset extremely challenging. Aside from the evaluation of quality for super-resolved images, we conducted further analysis to assess their performance in downstream biology tasks such as cell segmentation. Several methods achieved promising results compared to LR and HR images over cell detection/segmentation tasks. SR-CACO-2 can contribute to driving progress in the design of SISR models for microscopy. This extends from fixed-imaging to live-imaging. SISR models trained with our dataset can be employed for live-imaging videos captured with low quality under reduced light exposure to preserve cells. This leads to fast imaging which allows for the observation of instantaneous inter-cellular events with less damage to the cells. Accurately upscaling these videos with SISR models may ultimately open the door to performing biological tasks such as cell tracking, counting, and segmentation.
@@ Acknowledgments
Appendix A Overview of SR-CACO-2 Dataset
@@ A.1 Dataset statistics
We provide a nutrition label for our SR-CACO-2 dataset (Fig.14). Furthermore, Tab.6 presents more statistics about ROI distribution per HR tile for SR-CACO-2 dataset. Note that cells cover only a small part of the tile in general.
@@ A.2 Data Format and Organization
where ID $\in$ [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, ’7’, ’8’, ’9’, ’10’, ’11’, ’12’, ’13’, ’14’, ’15’, ’16’, ’17’, ’18’, ’19’, ’20’, ’21’, ’22’]. Note that a tile contains all three markers, one in each plane: plane 0 for CELL0, plane 1 for CELL1, a plane 2 for CELL2. The names of the patches follow this format: